OpenStudy (anonymous):

Suppose y varies directly with x, and y =14 when x = -4. What is the value of y when x = -6

OpenStudy (anonymous):

y varies directly with x means that y = cx lets figure out c by plugging in the x y combination that we know y = cx 14 = c*(-4) Solve that for c. What do you get?

OpenStudy (anonymous):

Using that formula I get 10

OpenStudy (anonymous):

14 = c*(-4) Divide both sides by -4, you get c = -14/4 or -7/2 so....y = (-7/2)x is our formula Now plug in -6 for x in our new formula to get y when x = -6

OpenStudy (anonymous):

(-7/2)(-6)=42/2=21

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.